OpenStudy (anonymous):

this is my last question. An equation that defines y as a function of x is given. Solve for y in terms of x, and replace y with the function notation f(x). 9x2 + 7y = 6

OpenStudy (anonymous):

the way its written with "9x2" doesnt make sense

OpenStudy (anonymous):

9x^2

jimthompson5910 (jim_thompson5910):

\[\large 9x^2 + 7y = 6\] \[\large 7y = 6-9x^2\] \[\large 7y = -9x^2+6\] \[\large y = \frac{-9x^2+6}{7}\] \[\large f(x) = \frac{-9x^2+6}{7}\]